OpenStudy (anonymous):

r(x)=(1/x-5)+2 write this as a rational expression by adding the terms

OpenStudy (campbell_st):

well you need a common denominator or (x - 5) so you need to simplify \[r(x) = \frac{1}{x -5} + \frac{2}{1} \times \frac{x - 5}{x -5}\] hope this helps.

OpenStudy (anonymous):

do I just multiply them then?

OpenStudy (campbell_st):

yep... so its \[r(x) = \frac{1}{x -5} + \frac{2 \times (x -5)}{x - 5}\] simplify the 2nd fraction... and then just add the numerators

OpenStudy (anonymous):

2x-10/x-5

OpenStudy (campbell_st):

thats the 2nd fraction... so you have \[r(x) = \frac{ 1 + 2x -10}{x - 5}\] just simplify it 1 - 10 =

OpenStudy (anonymous):

-9

OpenStudy (campbell_st):

yep so the answer is just \[r(x) = \frac{2x -9}{x - 5}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

so to determine the x intercept of this would you just set it equal to zero and solve? I got x=9/2

OpenStudy (stamp):



OpenStudy (stamp):

what value of x provides y = 0 solve 2x - 9 = 0 2x = 9 x = 9/2 x-int at (9/2, 0)

OpenStudy (anonymous):

and the y intercept is 9/5, 0

OpenStudy (stamp):

when x =0, what is your y value to solve y-int (0, y) f(0) = 9/5 y-int at (0, 9/5)